The management of Firebolt Industries Inc. manufactures gasoline and diesel engines through two production departments, Fabrication and Assembly. Management needs accurate product cost information in order to guide product strategy. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering the multiple production department factory overhead rate method. The following factory overhead was budgeted for Firebolt:

1

Fabrication Department factory overhead

$614,800.00

2

Assembly Department factory overhead

246,750.00

3

Total

$861,550.00

Direct labor hours were estimated as follows:

Fabrication Department 5,300 hours
Assembly Department 5,250
Total 10,550 hours

In addition, the direct labor hours (dlh) used to produce a unit of each product in each department were determined from engineering records, as follows:

Production Departments Gasoline Engine Diesel Engine
Fabrication Department 2.9 dlh 1.8 dlh
Assembly Department 1.8 2.9
Direct labor hours per unit 4.7 dlh 4.7 dlh
Required:
a. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the single plantwide factory overhead rate method, using direct labor hours as the activity base.*
b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department.*
c. (1) Recommend to management a product costing approach, based on your analyses in (a) and (b). (2) Give a reason for your answer.
*If required, round all per-unit answers to the nearest cent.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Fire bolt Industries Inc.
Answer a
Calculation of Plant wide OH Rate Fabricating Assembly Total Note
Manufacturing overhead             614,800.00         246,750.00 861,550.00 A
Direct Labor                 5,300.00             5,250.00     10,550.00 B
Predetermined OH Rate             81.66 C=A/B
Allocated OH to products Gasoline Engine Diesel Engine
Direct Labor                         4.70                     4.70 D
Predetermined OH Rate                       81.66                   81.66 See C
Allocated OH per unit                     384.00                 384.00 E=C*D
Answer b
Calculation of Departmental OH Rate Fabricating Assembly Note
Manufacturing overhead             614,800.00         246,750.00 F
Direct Labor                 5,300.00             5,250.00 G
Predetermined OH Rate                     116.00                   47.00 H=F/G
Allocated OH to products Gasoline Engine Diesel Engine
Direct Labor- Fabricating                         2.90                     1.80 I
Predetermined OH Rate                     116.00                 116.00 See H
OH per unit                     336.40                 208.80
Direct Labor- Assembly                         1.80                     2.90
Predetermined OH Rate                       47.00                   47.00
OH per unit                       84.60                 136.30
Total Allocated OH per unit                     421.00                 345.00 J=H*I
Answer c 1
(1) Recommend to management a product costing approach, based on your analyses in (a) and (b).
Management should use Departmental wide OH Rate.
Answer c 2- Reason
Because it shows more correct allocation and cost per unit of products on the basis of usage of direct labor hours in both fabricating and assembly department.